OpenStudy (anonymous):

Find the derivative of f(x) = 6/x at x = -2.

OpenStudy (vijay):

-3/2

OpenStudy (anonymous):

Can you explain how you got that cuz i wanna be able to do the other problems i have

OpenStudy (anonymous):

Do you just substitute?

OpenStudy (vijay):

differentiate then substitute

OpenStudy (anonymous):

differentiate?

OpenStudy (vijay):

yes, derivative means u have to differentiate

OpenStudy (anonymous):

How do you differentiate?

OpenStudy (vijay):

f(x) = 6 x^(-1) df/dx = (-1)* 6 *X^(-2)

OpenStudy (vijay):

now substitute x= -2

OpenStudy (vijay):

which is give -3/2

OpenStudy (anonymous):

sorry to keep saying thisbut i'm still confused. as tohow you got this "df/dx = (-1)* 6 *X^(-2)"

OpenStudy (vijay):

f(x) = 6 * x^(-1)

OpenStudy (vijay):

now if u do the differentiation

OpenStudy (agent0smith):

Use the power rule to differentiate... multiply by the power, and reduce the power by one, or: \[\Large \frac{ d }{ dx } x^n = n x ^{n-1}\] \[ \Large f(x) =\frac{ 6 }{ x} = 6x^{-1}\] So multiply by the power (-1) and reduce the power by one (-1 -1 = -2):\[\Large \frac{ d }{ dx} (6x^{-1}) = -1*6 x^{-2}\]

OpenStudy (ybarrap):

Following @agent0smith, for this problem, n = 1 and x is in the denominator: \[6*\frac{ d}{ dx } (1/x^{n}) = 6\frac{ d}{ dx } x^{-n} = -n *6 x^{-n -1}=6\frac{-n}{x^{n+1}}\]

OpenStudy (anonymous):

Thank you^^